the sum of peters age and his fathers age is 40 years. in two years time peters father will be 3 times as old as peter. how old is peter and his father?
work out the answer by using simultaneous equation method.
thanks

Answer by checkley79(3341)

P+D=40
3(P+2)=(D+2)
3P+6=D+2
3P-D=2-6
3P-D=-4
P+D=40 ADD.
-------------
4P=36
P=36/4
P=9 AGE OF PETER NOW.
9+D=40
D=40-9
D=31 AGE OF DAD NOW..
PROOF:
3(9+2)=31+2
3*11=33
33=33